# Why do zeroes happen? A model-based approach for demand classification

### v1

@@ 2.4 Stockout/Out-of-Stock identification
Generally, the retail operations literature has focused on the customer reactions to an Out-of-Stock (OOS) situation from the marketing perspective (e.g. Campo et al., 2000; Verbeke et al., 1998), while some studies look at the extent and root cause analysis of stockout situations, mainly connecting these to either retail store replenishment causes or upstream problems (Aastrup and Kotzab, 2010). There seem to be two main methods for auditing OOS/stockout situations in practice using: (1) shelves images/scanning (e.g. Šikić et al., 2024; Rosado et al., 2016) or (2) a data-driven approach based on point-of-sales (POS) data. Although both methods might be expensive and require additional tools and understanding, we would argue that the latter is easier to implement for most cases.
Fisher and Raman (2010) advocate for a data-driven analytical approach to improve retail supply chain performance by leveraging customer transaction data, demand forecasting, and inventory optimization techniques. They proposed to use dynamic inventory management, where retailers adjust stock levels based on demand patterns, seasonal fluctuations, profits and store-specific trends. Clearly, there is a need for automatic or semi-automatic detection of stockout via any available methods.
Papakiriakopoulos et al. (2009) proposed a rule-based decision support for the detection of OOS products based on heuristic rules. Their method analyses POS data, inventory records, and historical sales trends, applying predefined rules to identify anomalies – such as sudden sales drops or discrepancies between stock levels and sales activity, that may indicate stockouts. Using an iterative process of physical audit and classification models with internal and external validation, the authors were able to detect about one third of the OOS cases accurately. The authors note that while the system demonstrates acceptable levels of predictive accuracy and problem coverage, it may not account for all variables influencing shelf stock levels, such as sudden changes in consumer behaviour or external factors affecting sales patterns. Chuang (2018) expanded these ideas by including cost factors in the modelling.
Finally, Fildes et al. (2022) presented forecasting research in presence of stockouts in retail setting. The stockouts themselves can be complete or partial. In case of the former, the product is unavailable and we record zero sales. The latter implies that we run out of product in the middle of the day, so we cannot satisfy the whole demand. In this paper, we focus only on the complete stockouts, because the partial ones can only be identified if the stock system records the data correctly – it is not always possible to identify the partial stockouts correctly just by analysing sales.
@@ 3 Model-based Identification Method
The demand with naturally occurring zeroes can be considered as a proper intermittent. This type of demand can be count or fractional as well, similarly to the regular one. Depending on whether it is a regular or intermittent time series, an analyst can use an appropriate forecasting technique. For instance, if the demand is “count regular”, a Negative Binomial based model can be used. If it is “fractional regular”, any conventional forecasting model, such as ETS or ARIMA, or any machine learning technique, can be used. Following Syntetos and Boylan (2005), we propose to split intermittent intro smooth and lumpy to capture different customers behaviour.
@@ 3.1 Identifying stockouts
To achieve this, we extract the demand intervals $q_{j_{t}}$ from the data, similar to how they were originally proposed by Croston (1972), by calculating the difference between the indices of consecutive non-zero observations. This means that two consecutive non-zero demands will have a demand interval of one between them. Assuming that demand occurrence follows a Bernoulli distribution with a time-varying probability $p_{t}$, the demand intervals should follow the Geometric distribution:
@@
To estimate the time-varying probability, we fit a Friedman’s Super Smoother (Friedman, 1984) to the series $q_{j_{t}}$, accounting for potential changes in occurrence probability, for example due to demand becoming obsolete. While other smoothers could be used instead of this one (e.g. LOWESS by Cleveland, 1979), we found that the Super Smoother is sensitive enough to capture the potential changes in the demand intervals length. The smoothed series $\hat{q}_{j_{t}}$ is then used to compute $\hat{p}_{t}=\frac{1}{\hat{q}_{j_{t}}}$. Next, we identify observations exceeding a threshold $\nu$, determined by the quantile function of the Geometric distribution. For instance, setting $\nu$ to be equal to 0.99 marks the top 1% of values as potential stockouts.
To demonstrate the logic with stockouts identification, we consider an example of an intermittent time series (N10514 from the M5 dataset Makridakis et al., 2022), which is shown in Figure 2.
@@
(d) Count vs Fractional.
@@
Once stockouts are identified, they need to be addressed. The simplest approach is to remove them from the original data before proceeding to the next step in demand identification.
@@ 3.2 Automated Identification of Demand (AID)
If removing stockouts leaves no zeroes, the demand is regular, and we only need to determine if it is fractional or count. If zeroes remain, the demand is intermittent, requiring further classification as lumpy or smooth, since lumpy demand necessitates a separate model for demand occurrence.
To do the checks for the type of intermittent demand, we propose the following procedure. First, we fit a smooth line (such as Friedman’s Super Smoother or LOWESS, respectively by Friedman, 1984; Cleveland, 1979) to the overall demand $y_{t}$, the demand sizes $z_{t}$ and to the demand occurrence $o_{t}$, capturing the potential changes in the dynamics of the data. We thus obtain three smoothed series, $\hat{y}_{t}$, $\hat{z}_{t}$ and $\hat{p}_{t}$ respectively. After that, we use them in fitting several regression models for each of the categories of demand:
@@
    Regular Fractional – the model applied to the data itself, $y_{t}\sim\mathcal{N}(\beta_{0}+\beta_{1}\hat{y}_{t},\sigma_{y}^{2})$, where $\beta_{j}$ is a parameter of the model;
@@
    Regular Count – Negative Binomial distribution, $y_{t}\sim\mathcal{NB}(\beta_{0}+\beta_{1}\hat{y}_{t},s_{y})$, where $s_{y}$ is the scale of distribution, estimated together with other parameters of the model. We use this distribution as one of the most flexible count ones;
@@
    Lumpy Intermittent Fractional – the mixture distribution model: $y_{t}=o_{t}z_{t}$, where for demand sizes, $z_{t}\sim\mathcal{N}(\log\beta_{0}+\beta_{1}\hat{z}_{t},\sigma_{z}^{2})$ and for the probability of occurrence, $o_{t}\sim\mathrm{Bernoulli}(\beta_{0}+\beta_{1}\hat{p}_{t})$;
@@
Figure 3: Flowchart of the AID algorithm.
The logic in the application of the above models is as follows. In the first step, we check whether there are any zeroes left after removing stockouts to decide whether the data is regular or intermittent. If there are some left, they must be natural, meaning that the demand is indeed Intermittent. When this is decided, we then move to the next stage, calculating information criteria (such as AIC by Akaike, 1974) and comparing specific models to decide what specific type of demand we have. In case of the regular demand, we compare AIC of the model (1) with the one of the model (2), selecting the model, which has the lowest value. If the demand was identified as intermittent, and has fractional values, we then compare AIC of the models (3) with (4) to decide what specific type of demand we have: Smooth Intermittent or the Lumpy Intermittent. On the other hand, if the data has integer values only, it can be modelled using either a count data model or a fractional one. To determine, which one is better, we compare AIC of the models (3), (4), (5) and (6). For simplicity, the whole proposed algorithm is summarised in the flowchart in Figure 3.
@@ 4.2 Demand Identification
For this part of the experiment, we simulated data from six DGPs, each representing different types of demand. These are inspired by the iETS model of Svetunkov and Boylan (2023):
@@
The data was generated using the sim.es() function from the smooth package in R (Svetunkov, 2024b). The resulting series looked similar to the data shown in Figure 1. The simulation was done for the samples of 30, 60, 100, 400 and 1000 observations.
We then applied the aid() function with a confidence level of 0.999 for stockout detection and recorded how often each demand category was correctly identified. Using such a high confidence level reduces the likelihood of zeros being misclassified as stockouts, though the AID approach may still occasionally flag some incorrectly. Lowering the confidence level would result in more intermittent series being classified as regular, as more zeros would be treated as artificially occurring. Another option would be to set the level equal to one, thus switching off the stockouts detection mechanism completely, but we decided not to do that because we wanted to see the impact of the mechanism on the final classification.
@@
Overall, the algorithm struggled to correctly identify count data, particularly with smaller samples. This is because fractional demand models can often be applied effectively, even when the data is count-based (as shown, for example, by Svetunkov and Boylan, 2023). The algorithm also had difficulty identifying “Smooth Intermittent Fractional” demand, even with the largest sample of 1000 observations. This was mainly due to the AID approach frequently flagging zeros as stockouts for this data, causing the demand to be misclassified as “Regular Fractional”. In some cases, series in this group were also misclassified as “Lumpy Intermittent Fractional”, although this misclassification decreased with the increase of the sample size. This is likely because models for lumpy intermittent demand can also be effectively applied to smooth intermittent series.
Overall, the simulation experiment shows that the proposed approach performs well with larger samples but makes some mistakes with smaller ones. It effectively detects regular and lumpy intermittent demand, though it struggles with the smooth category due to limitations in the stockout detection algorithm. The distinction between count and fractional demand is not particularly pronounced from a modelling perspective, which leads the AID method to occasionally misclassifying the count ones. Nevertheless, the approach serves as a solid starting point for further analysis. To improve its accuracy, the confidence level for stockout detection should be carefully adjusted to prevent naturally occurring zeros from being flagged as stockouts. Additionally, if identifying count data is particularly important, it is advisable not to rely solely on automated detection but to directly verify whether the demand sizes are integer-valued.
@@ 5.2 Stockouts detection
We applied our classification scheme to the data via the aidCat() function from the greybox package in R (Svetunkov, 2024a) with level=0.999. It identified stockouts for each time series, the distribution of which is summarised in Figure 9.
@@ 5.4 Forecasting approaches and features
    LightGBM (Ke et al., 2017) – because of its speed and ability to handle large datasets like the one we had. We did our experiments using the lightgbm package in R (Shi et al., 2024);
@@
    Pooled regression applied to the whole dataset – to see whether the findings hold for a linear model that does not have as much flexibility as the LightGBM. This was done using the alm function from the greybox package in R (Svetunkov, 2024a);
@@
    SmoothSales – Smoothed original series. The smoothing was done using Friedman’s Super Smoother (Friedman, 1984) via the supsmu() function from the stats package in R (R Core Team, 2020). We used it because it is more sensitive to the local level changes than LOWESS. In cases of small samples (less than 7 non-zero observations), we substituted the values by the in-sample mean. For the holdout part of the data, we repeated the last available smoothed value in the sample for each series;
@@
  5. 9.
    Probability – the smoothed binary demand occurrence variable (the estimate of the probability of occurrence), done after removing the stockouts.
While there can be many other features that could be added to the experiment (such as ETS components or smoothed quantiles of the data), the aim was not to find the most suitable set of features, but rather to better understand the impact of our classification on forecast accuracy.
@@
    Conventional – one LightGBM approach applied directly to the full dataset ignoring the stockouts feature and using features (1), (2), (3), (4), and (6);
@@
    Category Partial – three LightGBMs, one applied to the data which was flagged as “Regular” in the manner similar to (2), and the other two applied to the data flagged as “Intermittent” in the manner similar to (3);
@@
    Category Full – Similar to (4), but with the split of the data into Regular/Smooth Intermittent/Lumpy Intermittent. The “Full” approach was applied to the Regular demand and two separate “Mixture” approaches were used for the Smooth and Lumpy intermittent demand. With this split, we want to see whether the more thorough split into categories brings any improvement;
@@
We should note at this stage that we could not come up with unique features that would support the split into smooth and lumpy intermittent categories. The separation into two has a reasonable theoretical rationale but does not yet has distinct characteristics.
@@
Finally, we only focused on measuring the point forecast accuracy of approaches, by calculating the Root Mean Squared Scaled Error (RMSSE) from Makridakis et al. (2022), originally motivated by Athanasopoulos and Kourentzes (2023).
5.5 Results of the experiment

### v2

@@ 2.4 Stockout/Out-of-Stock identification
We note that our paper does not focus on the censored demand and its implications to the inventory decisions, although there are many papers on the topic (see, for example, Lau1996; Sachs2014; Trapero2024). Instead, we focus on the demand classification, arguing that stockouts detection is one of the first important steps an analyst needs to make.
Generally, the retail operations literature has focused on the customer reactions to an Out-of-Stock (OOS) situation from the marketing perspective (e.g. Campo2000; Verbeke1998), while some studies look at the extent and root cause analysis of stockout situations, mainly connecting these to either retail store replenishment causes or upstream problems (Aastrup2010). There seem to be two main methods for auditing OOS/stockout situations in practice using: (1) shelves images/scanning (e.g. iki2024; Rosado2016) or (2) a data-driven approach based on point-of-sales (POS) data. Although both methods might be expensive and require additional tools and understanding, we would argue that the latter is easier to implement for most cases.
fisher2010 advocate for a data-driven analytical approach to improve retail supply chain performance by leveraging customer transaction data, demand forecasting, and inventory optimization techniques. They proposed to use dynamic inventory management, where retailers adjust stock levels based on demand patterns, seasonal fluctuations, profits and store-specific trends. Clearly, there is a need for automatic or semi-automatic detection of stockout via any available methods.
Papakiriakopoulos2009 proposed a rule-based decision support for the detection of OOS products based on heuristic rules. Their method analyses POS data, inventory records, and historical sales trends, applying predefined rules to identify anomalies – such as sudden sales drops or discrepancies between stock levels and sales activity, that may indicate stockouts. Using an iterative process of physical audit and classification models with internal and external validation, the authors were able to detect about one third of the OOS cases accurately. The authors note that while the system demonstrates acceptable levels of predictive accuracy and problem coverage, it may not account for all variables influencing shelf stock levels, such as sudden changes in consumer behaviour or external factors affecting sales patterns. Chuang2018 expanded these ideas by including cost factors in the modelling.
Finally, Fildes2022 presented forecasting research in presence of stockouts in retail setting. The stockouts themselves can be complete or partial. In case of the former, the product is unavailable and we record zero sales. The latter implies that we run out of product in the middle of the day, so we cannot satisfy the whole demand. In this paper, we focus only on the complete stockouts, because the partial ones can only be identified if the stock system records the data correctly – it is not always possible to identify the partial stockouts correctly just by analysing sales.
@@ 3 Model-based Identification Method
The demand with naturally occurring zeroes can be considered as a proper intermittent. This type of demand can be count or fractional as well, similarly to the regular one. Depending on whether it is a regular or intermittent time series, an analyst can use an appropriate forecasting technique. For instance, if the demand is “count regular”, a Negative Binomial based model can be used. If it is “fractional regular”, any conventional forecasting model, such as ETS or ARIMA, or any machine learning technique, can be used. Following Syntetos2005, we propose to split intermittent into smooth and lumpy to capture different customers behaviour, but our split comes from the understanding of the possible demand process rather than the minimum of MSE.
@@ 3.1 Identifying stockouts
To achieve this, we extract the demand intervals $q_{j_{t}}$ from the data, similar to how they were originally proposed by Croston1972, by calculating the difference between the indices of consecutive non-zero observations. This means that two consecutive non-zero demands will have a demand interval of one between them. Assuming that demand occurrence follows a Bernoulli distribution with a time-varying probability $p_{t}$, the demand intervals should follow the Geometric distribution:
@@
To estimate the time-varying probability, we fit a Friedman’s Super Smoother (friedman1984variable) to the series $q_{j_{t}}$, accounting for potential changes in occurrence probability, for example due to demand becoming obsolete. While other smoothers could be used instead of this one (e.g. LOWESS by Cleveland1979), we found that the Super Smoother is sensitive enough to capture the potential changes in the demand intervals length. The smoothed series $\hat{q}_{j_{t}}$ is then used to compute $\hat{p}_{t}=\frac{1}{\hat{q}_{j_{t}}}$. Next, we identify observations exceeding a threshold $\nu$, determined by the quantile function of the Geometric distribution. For instance, setting $\nu$ to be equal to 0.99 marks the top 1% of values as potential stockouts.
To demonstrate the logic with stockouts identification, we consider an example of an intermittent time series (N10514 from the M5 dataset Makridakis2022), which is shown in Figure 2.
@@
(d) Original series with detected stockouts.
@@
Once stockouts are identified, they need to be addressed. The simplest approach is to remove them from the original data before proceeding to the next step in demand identification. We note that we remove the stockout observations purely for the classification purposes to decide whether the demand is indeed intermittent or regular. But we do not recommend doing that for forecasting or inventory management.
@@ 3.2 Automated Identification of Demand (AID)
If removing stockouts leaves no zeroes, the demand is regular, and we only need to determine if it is fractional or count. If zeroes remain, the demand is intermittent, and we can determine whether it is lumpy or smooth. Thus AID can be considered as a two-level approach, where the first one (regular/intermittent) is determined automatically after the stockouts are removed, and the second one requires further investigation.
To do the checks for the type of intermittent demand, we propose the following procedure. First, we fit a smooth line (such as Friedman’s Super Smoother or LOWESS, respectively by friedman1984variable; Cleveland1979) to the overall demand $y_{t}$, the demand sizes $z_{t}$ and to the demand occurrence $o_{t}$, capturing the potential changes in the dynamics of the data. We thus obtain three smoothed series, $\hat{y}_{t}$, $\hat{z}_{t}$ and $\hat{p}_{t}$ respectively. After that, we use them in fitting several regression models for each of the categories of demand:
@@
[…]
@@
[…]
@@
[…]
@@
Figure 3: Flowchart of the AID algorithm. IC is the abbreviation of the term “Information Criterion”.
The logic in the application of the above models is as follows. In the first step, we check whether there are any zeroes left after removing stockouts to decide whether the data is regular or intermittent. If there are some left, they must be natural, meaning that the demand is indeed Intermittent. When this is decided, we then move to the next stage, calculating information criteria (such as AIC by Akaike1974) and comparing specific models to decide what specific type of demand we have. In case of the regular demand, we compare AIC of the model (1) with the one of the model (2), selecting the model, which has the lowest value. If the demand was identified as intermittent, and has fractional values, we then compare AIC of the models (3) with (4) to decide what specific type of demand we have: Smooth Intermittent or the Lumpy Intermittent. On the other hand, if the data has integer values only, it can be modelled using either a count data model or a fractional one. To determine, which one is better, we compare AIC of the models (3), (4), (5) and (6). For simplicity, the whole proposed algorithm is summarised in the flowchart in Figure 3.
We should also note that the AID algorithm can be extended by the inclusion of trends, seasonal components, and exogenous variables. However, we argue that the logic of the classification should not change in this case, and, for example, the model for the lumpy demand would be selected in both cases of having or not having the additional features. This is because all the models under comparison are compared conditional on the same information, and addition of features would reduce information criteria for all of them in a similar manner, thus not affecting the ranking of the models based on IC.
@@ 4.2 Demand Identification
For this part of the experiment, we simulated data from six DGPs, each representing different types of demand. These are inspired by the iETS model of Svetunkov2023b:
@@
The data was generated using the sim.es() function from the smooth package in R (SvetunkovSmooth). The resulting series looked similar to the data shown in Figure 1. The simulation was done for the samples of 30, 60, 100, 400 and 1000 observations.
We then applied the aid() function with a confidence level of 0.999 for stockout detection and recorded how often each demand category was correctly identified. Using such a high confidence level reduces the likelihood of zeroes being misclassified as stockouts, though the AID approach may still occasionally flag some incorrectly. Lowering the confidence level would result in more intermittent series being classified as regular, as more zeroes would be treated as artificially occurring. Another option would be to set the level equal to one, thus switching off the stockouts detection mechanism completely, but we decided not to do that because we wanted to see the impact of the mechanism on the final classification.
@@
Overall, the algorithm struggled to correctly identify count data, particularly with smaller samples. This is because fractional demand models can often be applied effectively, even when the data is count-based (as shown, for example, by Svetunkov2023b). The algorithm also had difficulty identifying “Smooth Intermittent Fractional” demand, even with the largest sample of 1000 observations. This was mainly due to the AID approach frequently flagging zeroes as stockouts for this data, causing the demand to be misclassified as “Regular Fractional”. In some cases, series in this group were also misclassified as “Lumpy Intermittent Fractional”, although this misclassification decreased with the increase of the sample size. This is likely because models for lumpy intermittent demand can also be effectively applied to smooth intermittent series.
In addition to this, we also conducted an experiment where outliers are added to the data, reflecting the situation of promotions that are not taken into account by the AID algorithm. They were added randomly in the sample when non-zero values occurred by multiplying those values by two. The idea was to see how the results would change if we have some elements of the structure that are ignored by AID. The results of this additional experiment are shown in Figure 9.
Figure 9: Demand identification for each of categories of DGPs. The case of missed promotions.
Comparing Figures 8 and 9, we can see that the only thing that changes in this situation is the accuracy of the identification of the fractional vs count demand on small samples: the demand in those two categories is now identified better than in the case of no promotions. This is probably because the multiplication by two leads to values that look more apparent as fractional/count.
In summary, the simulation experiment shows that the proposed approach performs well with larger samples but makes some mistakes with smaller ones. It effectively detects regular and lumpy intermittent demand, though it struggles with the smooth category due to limitations in the stockout detection algorithm. The distinction between count and fractional demand is not particularly pronounced from a modelling perspective, which leads the AID method to occasionally misclassifying the count ones. Nevertheless, the approach serves as a solid starting point for further analysis. To improve its accuracy, the confidence level for stockout detection should be carefully adjusted to prevent naturally occurring zeroes from being flagged as stockouts. Additionally, if identifying count data is particularly important, it is advisable not to rely solely on automated detection but to directly verify whether the demand sizes are integer-valued.
@@ 5.2 Stockouts detection
We applied our classification scheme to the data via the aidCat() function from the greybox package in R (Svetunkovgreybox) with level=0.999. It identified stockouts for each time series, the distribution of which is summarised in Figure 10.
@@ 5.4 Forecasting approaches and features
    LightGBM (Ke2017) – because of its speed and ability to handle large datasets like the one we had. We did our experiments using the lightgbm package in R (ShiLightGBM);
@@
    Pooled regression applied to the whole dataset – to see whether the findings hold for a linear model that does not have as much flexibility as the LightGBM. This was done using the alm function from the greybox package in R (Svetunkovgreybox);
@@
    Category Partial – the demand categories according to AID, including only either “intermittent”, or “regular”. The idea of this feature is to see whether the split to the these two categories would improve the performance of models;
@@
    Category Full – the three demand categories according to AID, including “regular”, “smooth intermittent” and “lumpy intermittent”. Similarly to the previous feature, this one explores whether the further split brings value;
  4. 8.
    SmoothSales – Smoothed original series. The smoothing was done using Friedman’s Super Smoother (friedman1984variable) via the supsmu() function from the stats package in R (RCore2020). We used it because it is more sensitive to the local level changes than LOWESS. If the smoothed line was below zero, we would substitute its values with zero, reflecting the fact that the sales cannot be negative. In cases of small samples (less than 7 non-zero observations), we substituted the values by the in-sample mean. For the holdout part of the data, we repeated the last available smoothed value in the sample for each series;
  5. 9.
@@
  7. 11.
    Probability – the smoothed binary demand occurrence variable (the estimate of the probability of occurrence), done after removing the stockouts. If the smoothed line was above one or below zero, we would substitute its values with the boundary values;
  8. 12.
    Seasonality – sine and cosine curves with periodicity of 52 (i.e. $\sin\left(\frac{2\pi j}{52}\right)$) were added as well. These help capturing the week-of-year seasonality by non-linear machine learning approaches (such as LightGBM), but are not useful for linear ones, such as regression (Barrow2015).
We remark that we do not add the feature of count/fractional demand because our experiments showed that while it is theoretically appealing, it does not bring any value in terms of forecasting accuracy. Furthermore, while there can be many other features that could be added to the experiment (such as ETS components or smoothed quantiles of the data), the aim was not to find the most suitable set of features, but rather to better understand the impact of our classification on forecast accuracy.
@@
    Conventional – one LightGBM approach applied directly to the full dataset ignoring the stockouts feature and using features (1), (2), (3), (4), (8) and (12);
@@
    Category Partial – Same as the Mixture approach, but with the feature 6 added to both demand occurrence and the demand sizes parts of the approach;
@@
    Category Full – Same as the Category Partial approach, but with the feature 7 instead of 6.
@@
We should note that in this section, we only focused on measuring the point forecast accuracy of approaches, by calculating the Root Mean Squared Scaled Error (RMSSE) from Makridakis2022, originally motivated by Athanasopoulos2020. But an important aspect in the evaluation is that it needs to be done on the test set values that do not have stockouts. If the stockouts are not removed from the test set, the evaluation would point towards the models that are the most accurate at predicting sales instead of demand. We want to avoid that, so we dropped series that had only zeroes in the test set.
Finally, we conducted additional experiment using the count/fractional as a feature in the models, but the performance of LightGBM and regression was similar to the “Category Partial” of the respective models. We argue that the fact that the demand is either fractional or count does not matter that much for the forecasting accuracy, but it might be important for inventory decisions if we deal with a mixture of products.
5.5 Forecasting results